OpenStudy (anonymous):

3x2 - 4x = -2 by completing the square.

OpenStudy (anonymous):

3x^2 - 4x + 2 = 0 3x^2 - 5x + 2 = -x 3x^2 - 3x -2x + 2 = -x 3x(x-1) - 2(x-1) = -x (3x-2)(x-1) = -x